# NV-Embed: Improved Techniques for Training LLMs as Generalist Embedding Models

### v1

@@ 5.1 Experimental Details
We train Mistral 7B LLM model end-to-end with a contrastive loss using LoRA with rank 16, alpha 32 and dropout rate of 0.1. We use Adam optimizer with 500 warm-up steps and learning rate 2e-5 for first stage and 1.5e-5 for second stage with linear decay. The model is finetuned with 128 batch size, where each batch is composed of a query paired with 1 positive and 7 hard negative documents. We train using Bfloat16, and set the maximum sequence length as 512 tokens. The special <BOS> and <EOS> tokens are appended at the start and end of given query and documents. The whole training is conducted in two stages where the model is initially trained on retrieval datasets utilizing in-batch negative technique. Subsequently, the model is trained with blended datasets with both retrieval and non-retrieval embedding tasks.
@@ 5.2 MTEB Results
We evaluate the proposed NV-Embed model on the full MTEB benchmark (Muennighoff et al., 2022) encompassing 15 retrieval datasets, 4 reranking datasets, 12 classification datasets, 11 clustering datasets, 3 pair classification datasets, 10 semantic textual similarity datasets, and 1 summarization dataset.
Table 1 shows the averaged MTEB scores for overall performance and seven sub-category tasks compared to all frontier models on the MTEB leaderboard [^2] https://huggingface.co/spaces/mteb/leaderboard. Our NV-Embed model achieves a new record high score of 69.32 on the MTEB benchmark with 56 tasks and also attains the highest score of 59.36 on 15 retrieval tasks originally from the BEIR benchmark (Thakur et al., 2021).
Based on quantitative leaderboard results, we compare our NV-Embed with the recent frontier embedding models. The e5-mistral-7b-instruct (Wang et al., 2023b) and google-gecko (Lee et al., 2024a) utilize proprietary synthetic data to train their model in a single stage manner. In contrast, we recognize that retrieval task presents greater difficulty compared to the other embedding tasks and prioritizes our training strategy on fine-tuning the model for retrieval first, followed by blending the remaining sub-tasks into instruction-tuning, leading to substantially improved BEIR and overall METB results.
SFR-Embedding (Meng et al., 2024) demonstrates competitive scores on the MTEB (67.56) and BEIR (59.0) benchmarks by continuing to finetune the e5-mistral-7b-instruct model (Wang et al., 2023b). However, it remains largely constrained by the architectural limitations of its parent model, such as the causal attention mask and the last token pooling method. In contrast, our NV-Embed model is trained starting from the Mistral 7B LLM rather than finetuning e5-mistral-7b-instruct. It features a new architecture that removes the unnecessary causal attention mask and further improves the sequence pooling mechanism with a latent attention layer. Table 4 provides a detailed summary of task-wise BEIR and MTEB benchmarks.
Table 2: Averaged MTEB scores on seven tasks after first stage training
@@
Table 3: Averaged MTEB scores on seven tasks after second stage training
@@
Table 4: Full BEIR and MTEB benchmark
Model Name E5-mistral-7b SFR-Embedding Voyage-large2-instruct EOS Mean Latent-attention Self-attention
bidirect causal bidirect causal bidirect causal bidirect causal
ArguAna 61.88 67.27 64.06 67.06 62.51 63.83 64.14 68.21 64.57 65.56 62.82
ClimateFEVER 38.40 36.41 32.65 33.92 31.05 34.09 31.52 34.72 33.38 34.78 33.05
CQADupStack 42.97 46.54 46.60 48.39 46.14 48.69 47.13 50.51 47.44 50.19 46.79
Dbpedia 48.90 49.06 46.03 48.03 46.28 49.12 48.17 48.29 47.46 47.74 47.96
FEVER 87.80 89.35 91.47 87.11 85.80 87.90 87.46 87.77 87.66 86.89 87.45
FiQA2018 56.62 60.55 59.76 59.72 56.86 60.84 57.73 63.10 60.01 62.81 58.83
HotpotQA 75.70 77.02 70.86 78.40 75.75 78.65 77.90 79.92 78.37 79.33 77.73
MSMARCO 43.10 43.41 40.60 46.10 45.60 46.23 45.80 46.49 46.10 46.80 46.00
NFCorpus 38.59 42.02 40.32 38.48 36.89 39.13 38.31 38.04 38.48 37.68 37.58
Natural Question 63.50 69.92 65.92 70.07 68.75 71.23 69.43 71.22 70.77 71.33 70.22
QuoraRetrieval 89.62 89.81 87.40 88.88 88.57 88.75 88.73 89.21 88.71 88.85 88.76
SCIDOCS 16.27 19.91 24.32 19.81 16.81 21.08 20.38 20.19 18.86 20.86 18.20
SciFact 76.41 78.06 79.99 77.21 75.48 77.53 78.22 78.43 79.17 76.83 77.38
TREC-COVID 87.33 87.10 85.07 85.34 84.42 85.87 84.44 85.88 85.81 83.75 84.64
Touche2020 26.39 29.00 39.16 27.36 28.00 27.66 28.84 28.38 28.08 26.21 28.26
BIOSSES 85.58 86.07 89.12 86.44 83.04 86.19 83.27 85.59 83.37 85.26 82.71
SICK-R 82.64 82.92 83.16 78.59 77.65 82.87 81.06 82.80 81.44 83.21 80.87
STS12 79.65 79.47 76.15 73.30 72.77 74.82 73.54 76.22 75.03 77.10 74.67
STS13 88.43 89.15 88.49 81.84 83.10 85.81 86.48 86.30 85.44 84.54 84.07
STS14 84.54 84.93 86.49 77.84 77.18 81.45 80.73 82.09 80.51 80.29 80.05
STS15 90.42 90.74 91.13 85.50 83.97 87.20 86.69 87.24 86.35 87.16 87.03
STS16 87.68 87.82 85.68 81.99 81.90 84.62 84.42 84.77 84.70 83.76 84.02
STS17 91.75 92.02 90.06 77.89 81.85 88.53 86.67 87.42 87.64 86.69 86.16
STS22 67.28 68.36 66.32 71.14 68.81 68.69 70.06 69.85 70.02 69.91 70.59
STSBenchmark 88.60 89.00 89.22 80.95 80.89 85.15 84.50 86.14 84.94 85.86 85.07
SummEval 31.40 31.16 30.84 30.36 29.12 30.49 31.82 31.20 31.87 30.11 31.40
SprintDuplicateQuestions 95.66 96.31 94.50 94.89 91.46 95.39 94.71 95.94 95.15 95.98 95.12
TwitterSemEval2015 81.62 81.52 86.32 70.94 65.73 75.81 67.80 78.73 69.06 76.31 72.41
TwitterURLCorpus 87.75 87.78 86.90 86.34 85.79 86.36 86.15 86.05 86.12 86.07 85.78
AmazonCounterfactual 78.69 77.93 77.60 94.69 93.87 94.48 94.10 95.12 93.88 94.78 93.64
AmazonPolarity 95.91 95.97 96.58 97.05 96.34 96.92 96.66 97.14 97.08 97.27 97.02
AmazonReviews 55.79 54.35 50.77 53.37 56.09 55.68 55.99 55.47 56.59 55.47 54.81
Banking77 88.23 88.81 86.96 87.93 87.08 89.13 88.55 90.34 89.08 89.76 89.37
Emotion 49.77 50.24 59.81 91.19 91.39 91.01 91.52 91.71 91.54 91.97 91.13
Imdb 94.78 94.79 96.13 97.15 96.34 96.83 96.26 97.06 96.69 97.13 97.03
MassiveIntent 80.57 79.99 81.08 78.89 76.37 80.13 78.91 80.07 80.42 79.88 79.31
MassiveScenario 82.39 82.20 87.95 81.53 78.79 81.80 81.94 81.74 83.24 81.92 81.96
MTOPDomain 96.12 96.36 98.86 96.28 95.34 96.58 96.08 96.51 95.93 96.98 96.33
MTOPIntent 86.11 86.30 86.97 88.16 85.58 88.58 88.19 89.77 88.60 88.97 89.27
ToxicConversations 69.59 69.33 83.58 93.17 93.16 92.86 92.57 92.60 93.41 93.21 93.04
TweetSentimentExtraction 63.72 63.64 71.55 79.62 79.19 80.68 79.86 80.64 80.41 81.10 79.72
Arxiv-P2P 50.45 52.08 51.81 53.60 53.23 53.45 53.23 53.76 53.21 53.51 53.24
Arxiv-S2S 45.50 47.38 44.73 48.23 48.71 49.52 48.79 49.59 49.01 49.61 49.00
Biorxiv-P2P 43.53 43.94 46.07 47.02 45.53 46.97 47.09 48.15 47.56 48.71 47.87
Biorxiv-S2S 40.24 41.14 40.64 43.99 43.52 44.03 43.26 44.74 43.76 45.36 44.58
Medrxiv-P2P 38.19 40.03 42.94 38.64 38.42 38.30 37.90 39.24 38.34 38.88 38.34
Medrxiv-S2S 37.45 39.00 41.44 36.67 38.52 36.94 36.61 36.98 36.88 37.53 36.98
Reddit 57.71 59.90 68.50 62.42 60.88 64.62 62.99 63.20 62.39 64.77 61.33
Reddit-P2P 66.49 67.64 64.86 67.67 63.92 68.01 66.59 68.01 66.85 68.17 65.99
StackExchange 73.10 74.25 74.16 72.44 68.49 77.07 73.35 74.99 72.36 76.58 72.03
StackExchange-P2P 45.91 46.78 45.10 41.32 36.40 40.95 37.79 42.04 38.99 41.87 38.28
TwentyNewsgroups 54.31 56.27 66.62 53.68 50.34 60.97 59.81 60.13 59.33 61.72 58.98
AskUbuntuDupQuestions 66.98 67.58 64.92 66.78 65.53 67.85 68.10 67.50 68.17 67.01 68.24
MindSmallRerank 32.60 32.72 30.97 30.53 28.71 31.10 30.98 30.82 31.50 31.81 31.63
SciDocsRR 86.33 86.58 89.34 88.05 86.33 87.71 87.24 87.26 87.24 87.38 86.74
StackOverflowDupQuestions 54.91 55.68 55.11 55.66 54.06 56.43 54.75 56.58 55.38 56.62 55.54
BEIR Average (15) 56.90 59.03 58.28 58.39 56.59 58.71 57.88 59.36 58.33 58.64 57.71
MTEB Average (56) 66.63 67.56 68.28 67.85 66.50 68.97 68.13 69.32 68.47 69.10 68.16
@@ 5.3 Ablation Study
We perform ablation studies to compare causal and bidirectional attention for contrastive trainng. We also compare the the proposed latent attention layer with other pooling methods.
@@ 5.3.2 Pooling Methods
To enhance performance beyond mean pooling, we experimented with adding the proposed latent-attention or self-attention layer (both followed by MLP) before mean pooling to address the issue of important information from key phrases being diluted. According to Table 2 and 3, self-attention does not provide additional accuracy improvements for the embedding capabilities of decoder-only LLMs (i.e., mean pooling 68.97 vs. self-attention 69.10 on 56 MTEB tasks). It even slightly reduces accuracy on 15 retrieval tasks (i.e., mean pooling 58.71 vs. self-attention 58.64). This is not surprising, as the LLM already has many self-attention layers to learn the representation, and adding an additional one does not bring significant additive value.
In contrast, the latent-attention layer proved beneficial for retrieval, classification, and STS subtasks, as shown in Table 3. Specifically, the nDCG@10 accuracy of the more challenging 15 retrieval tasks improved (i.e., mean pooling 58.71 vs. latent-attention 59.36). We hypothesize that this is due to the "dictionary learning" provided by the latent array, which offers more expressive representation. The latent-attention layer effectively learns output embedding representations from decoder-only LLMs, mitigating the information dilution caused by averaging the output embeddings.
@@ 6 Conclusion
In this work, we introduce NV-Embed model which presents novel architectural design and two-staged training procedure to substantially enhance the LLM capability as a generalist embedding model. For model architecture, we propose a latent attention layer to obtain expressive pooled embeddings and remove the unnecessary causal attention mask of decoder-only LLMs. For model training, we introduce a two-stage contrastive instruction-tuning scheme to sequentially improve the embedding tasks encompassing retrieval, classification, clustering, and semantic textual similarity. As of May 24, 2024, our NV-Embed model obtains a new record high score on the Massive Text Embedding Benchmark (MTEB) with 56 tasks and also attains the highest score on BEIR benchmark (15 retrieval tasks in the MTEB benchmark). Notably, we obtain state-of-the-art results using only publicly available data, without any synthetic data from frontier proprietary LLMs, such as GPT-4.
Appendix A Implementation Details

### v2

@@ 5.1 Experimental Details
We train Mistral 7B LLM model end-to-end with a contrastive loss using LoRA with rank 16, alpha 32 and dropout rate of 0.1. We use Adam optimizer with 50 warm-up steps and learning rate 2e-5 for first stage and 1.5e-5 for second stage with linear decay. The model is finetuned with 128 batch size, where each batch is composed of a query paired with 1 positive and 7 hard negative documents. Training samples from different datasets in Table 9 are uniformly sampled. We train using Bfloat16, and set the maximum sequence length as 512 tokens. The special <BOS> and <EOS> tokens are appended at the start and end of given query and documents. The whole training is conducted in two stages where the model is initially trained on retrieval datasets utilizing in-batch negative technique. Subsequently, the model is trained with blended datasets with both retrieval and non-retrieval embedding tasks.
@@ 5.2 MTEB Results
We evaluate the proposed NV-Embed model on the full MTEB benchmark (Muennighoff et al., 2022) across 56 tasks. Table 1 summarizes the averaged MTEB scores for seven sub-category tasks compared to the frontier models on the MTEB leaderboard [^2] https://github.com/embeddings-benchmark/mteb. Our initial model, namely NV-Embed-v1 get the score of 69.32 and obtain the No.1 position on the MTEB as of May 24, 2024 (detailed benchmark scores available in Table 2). We then further improve the model through the curation of training dataset, including adding more retrieval datasets, applying positive-aware hard-negative mining technique, using synthetic data generation process and constructing example-based multi-class labels. As a result, our NV-Embed-v2 model sets a new record high score of 72.31 and reclaimed No.1 (as of Aug 30, 2024) on the highly competitive MTEB leaderboard, further highlighting the sustained effectiveness of the proposed methods. In following sub-section 5.3, we will present the ablation studies on the design choices regarding the model architecture, training algorithm and the curation of training data.
Based on quantitative leaderboard results, we compare our NV-Embed with the recent frontier embedding models. The e5-mistral-7b-instruct (Wang et al., 2023b) and google-gecko (Lee et al., 2024a) utilize proprietary synthetic data to train their model in a single stage manner. In contrast, we recognize that retrieval task presents greater difficulty compared to the other embedding tasks and prioritizes our training strategy on fine-tuning the model for retrieval first, followed by blending the remaining sub-tasks into instruction-tuning, leading to substantially improved BEIR and overall MTEB results.
SFR-Embedding-2R (Meng et al., 2024b) demonstrates competitive scores on the MTEB (70.31) and BEIR (60.18) benchmarks by continuing to finetune the e5-mistral-7b-instruct model (Wang et al., 2023b). However, it remains largely constrained by the architectural limitations of its parent model, such as the causal attention mask and the last token pooling method. In contrast, our NV-Embed model is trained starting from the Mistral 7B LLM (Jiang et al., 2023) rather than finetuning e5-mistral-7b-instruct (Wang et al., 2023b). It features a new architecture that removes the unnecessary causal attention mask and further improves the sequence pooling mechanism with a latent attention layer. Table 3 and 11 provides a detailed scores of BEIR and MTEB benchmarks.
Table 2: Averaged MTEB scores on seven tasks after first and second stage training using only the publically available data and before applying the positive-aware hardnegative mining, synthetic data and example-based multi-class labeling. The averaged score 69.32 corresponds to NV-Embed-v1.
First stage training
@@
Second stage training
@@
Table 3: Averaged MTEB scores on seven embedding tasks after two stage training after applying the positive-aware hardnegative mining, synthetic data and example-based multi-class labeling. Note, the averaged score 72.31 corresponds to NV-Embed-v2.
Pool Type EOS Mean Latent-attention Self-attention
Retrieval (15) 62.13 60.30 61.81 61.01 62.65 61.15 61.17 60.53
Rerank (4) 60.02 59.13 60.65 59.10 60.65 59.36 60.67 59.67
Clustering (11) 58.24 57.11 57.44 57.34 58.46 57.80 58.24 57.11
PairClass. (3) 87.69 85.05 87.35 87.35 88.67 87.22 87.69 85.05
Classification (12) 90.10 90.01 89.49 89.85 90.37 90.49 90.10 90.01
STS (10) 82.27 81.65 84.35 84.35 84.31 84.13 84.22 83.81
Summar. (1) 30.25 32.75 30.75 30.88 30.70 30.90 30.93 31.36
Average (56) 71.63 70.85 71.71 71.38 72.31 71.61 71.61 70.6
@@ 5.3 Ablation Study
We conduct ablation studies to compare several training, architecture and data curation design choices: two-stage training, bidirectional attention, latent-attention pooling method, synthetic data and example-based multi-class labeling.
Table 4: Averaged MTEB scores on ablation studies for NV-Embed-v2: two stage training, multi-class data labeling, positive-aware hardnegative mining and synthetically generated dataset. In the third part of the table, HN represents hardnegative mining technique, AD means adding public retrieval datasets and SD refers to adding synthetically generated data. In the fourth part of the table, we also include NV-Embed-v1, which omits HN, AD, and SD in stage-one training and uses a label-based approach in stage-two training.
Section 5.3.1 Two stage training
Embedding Task Retrieval Rerank Cluster. PairClass. Class. STS Summ. Avg.
Single Stage (Inbatch Enabled)
61.25 60.64 57.67 87.82 86.6 83.7 30.75 70.83
Single Stage (Inbatch Disabled)
61.37 60.81 58.31 88.3 90.2 84.5 30.96 71.94
Two Stage Training
62.65 60.65 58.46 88.67 90.37 84.31 30.70 72.31
Reversed Two Stage
61.91 60.98 58.22 88.59 90.26 83.07 31.28 71.85
Section 5.3.4 Multi-class Classification and Clustering Labels in stage-two training
Embedding Task Retrieval Rerank Cluster. PairClass. Class. STS Summ. Avg.
Label-based approach
62.40 59.7 53.04 88.04 89.17 84.25 30.77 70.82
Example-based approach
62.65 60.65 58.46 88.67 90.37 84.31 30.70 72.31
Section 5.3.5 Hard-negative mining and Synthetically Generated Dataset in stage-one training
Embedding Task Retrieval Rerank Cluster. PairClass. Class. STS Summ. Avg.
[S0] Without HN, Without AD, Without SD
59.22 59.85 57.95 85.79 90.71 81.98 29.87 70.73
[S1] With HN, Without AD, Without SD
61.52 59.80 58.01 88.56 90.31 84.26 30.36 71.83
[S2] With HN, With AD, Without SD
62.28 60.45 58.16 88.38 90.34 84.11 29.95 72.07
[S3] With HN, With AD, With SD
62.65 60.65 58.46 88.67 90.37 84.31 30.70 72.31
NV-Embed-v1
Label-based approach + [S0]
59.36 60.59 52.80 86.91 87.35 82.84 31.2 69.32
5.3.1 Two-stage training
We compare the two-stage and single-stage training with and without the use of the in-batch negative technique, as shown in Table 4. We observe that our proposed two-stage training surpasses single-stage training because it allows the use of beneficial in-batch negatives for retrieval tasks in the first stage, while disabling the in-batch technique for non-retrieval tasks in the second stage. In contrast, single-stage training with in-batch negatives leads to significantly lower MTEB performance, especially in the classification sub-task. This accuracy degradation occurs because many classification tasks involve few-class labels (such as binary labels like True/False), meaning that the inbatch negative labels in the batch can actually be the positive label. While single-stage training without in-batch negatives produces more comparable results (MTEB scores: 72.31 for two-stage training vs. 71.94 for single-stage without in-batch), two-stage training significantly outperforms in the retrieval sub-tasks (BEIR scores: 62.65 for two-stage training vs. 61.37 for single-stage without in-batch). It is worth highlighting here that the retrieval is considered the most crucial sub-category for the advancement of RAG technology across the MTEB embedding tasks.
Lastly, we explore another research question: what happens if the order of two-stage training is reversed? To examine this, we further finetune the Single Stage (Inbatch disabled) model using only the retrieval datasets with enabling the inbatch negative technique and present the MTEB results in Table 4. While the retrieval score increased from 61.37 to 61.91 after the reversed two-staged training, it remains lower than the retrieval score of 62.65 achieved with our proposed two-stage training method. Furthermore, the scores on other embedding tasks, such as Clustering and STS, declined compared to the Single Stage (Inbatch disabled) approach. Consequently, the overall MTEB score for Reversed Two Stage (score: 71.85) is lower than our proposed Two-Stage Training (score: 72.31) as well as the Single Stage with Inbatch disabled (score: 71.94).
@@ 5.3.3 Pooling Methods
To enhance performance beyond mean pooling, we experimented with adding the proposed latent-attention or self-attention layer (both followed by MLP) before mean pooling to address the issue of important information from key phrases being diluted. According to Tables 2, self-attention does not provide additional accuracy improvements for the embedding capabilities of decoder-only LLMs (i.e., mean pooling 68.97 vs. self-attention 69.10 on MTEB tasks). It even slightly reduces accuracy on 15 retrieval tasks (i.e., mean pooling 58.71 vs. self-attention 58.64). Table 3 also shows the similar trends of NV-Embed-v2. This is not surprising, as the LLM already has many self-attention layers to learn the representation, and adding an additional one does not bring significant additive value.
In contrast, the latent-attention layer proved beneficial for majority of embedding tasks, as shown in Table 2 and 3. Specifically, the nDCG@10 accuracy of the more challenging 15 retrieval tasks improved (i.e., mean pooling 61.82 vs. latent-attention 62.65) in Table 3. We hypothesize that this is due to the "dictionary learning" provided by the latent array, which offers more expressive representation. The latent-attention layer effectively learns output embedding representations from decoder-only LLMs, mitigating the information dilution caused by averaging the output embeddings.
Table 5: Ablation study on using class/cluster labels vs. sampled class/cluster examples as positive and negative documents for multi-class classification and clustering tasks.
+/- Document Format Labels Examples
Emotion-Classification 90.83 93.38
MassiveIntent-Classification 84.94 86.10
MassiveScenario-Classification 90.18 92.17
MTOPDomain-Classification 98.84 99.25
MTOPIntent-Classification 88.55 94.37
Arxiv-Clustering-P2P 53.01 55.80
Arxiv-Clustering-S2S 49.19 51.26
Biorxiv-Clustering-P2P 45.38 54.09
Biorxiv-Clustering-S2S 42.67 49.60
Medrxiv-Clustering-P2P 37.58 46.09
Medrxiv-Clustering-S2S 36.82 44.86
Reddit-Clustering 59.83 71.10
Reddit-Clustering-P2P 72.58 74.94
StackExchange-Clustering 79.37 82.10
StackExchange-Clustering-P2P 48.59 48.36
TwentyNewsgroups-Clustering 58.41 64.82
Average (16) 64.80 69.27
5.3.4 Multi-class Classification and Clustering Labels
We compare the effect of using two possible techniques for constructing positive and negative documents for multi-class classification and clustering tasks. In label-based approach, the ground-truth class/cluster label corresponding to the example in the query is used as the positive document, and other class/cluster labels are sampled for negative documents. In example-based approach, another example from the same class/cluster as the example in the query is used as the positive document, and examples from other clusters are sampled for negative documents. We use random sampling to get a broad coverage across labels and examples. In this work, all 11 clustering datasets and 5 muti-class classification datasets are constructed as example-based approach. As shown in Table 4, the example-based approach leads to significant improvements over the label-based approach for both classification and clustering. Table 5 further shows the detailed ablation study of label-based and example-based labels for classification and clustering multi-class samples.
5.3.5 Hardnegative mining and Synthetically Generated Dataset
We provide a step-by-step curation of training dataset, incorporating the hard negative mining technique (S1), additional public retrieval data (S2), and synthetically generated data (S3). As shown in Table 4, the first step of adding the hard negative mining technique significantly boosted retrieval accuracy, with the BEIR score increasing from 59.22 to 61.52. In the next step (S2), we included more public retrieval datasets (HoVer, SciFact, Nfcorpus, MIRACL, Mr.Tydi) followed by synthetically generated data. Adding the public retrieval datasets further increased the retrieval score by 0.7 points. Finally, incorporating the synthetic dataset (S3) leads to a modest improvement in the overall MTEB scores, raising them by 0.24 points.
@@ 6 Conclusion
We introduced the NV-Embed model, a decoder-only LLM designed to outperform existing bidirectional models in general-purpose text embedding tasks. For model architecture, we propose a latent attention layer to obtain expressive pooled embeddings and remove the unnecessary causal attention mask of decoder-only LLMs. For training algorithm, we introduce a two-stage contrastive instruction-tuning scheme to sequentially improve the embedding tasks. By leveraging carefully curated datasets, hard-negative mining, synthetic data generation and example-based multi-class labeling, our approach achieve the superior accuracy across diverse embedding tasks. As a result, the series of NV-Embed models achieved and maintained the No.1 ranking on the MTEB leaderboard and also demonstrated superior accuracy in out-of-domain tasks in AIR Benchmark. The sustained effectiveness of NV-Embed highlights the importance of our proposed architecutre design, training procedures and dataset curation in achieving state-of-the-art performance in the evolving landscape of text embedding models.
7 Acknowledgment
We would like to extend our sincere gratitude to the NVIDIA Merlin team for their valuable collaboration and insightful discussions on building embedding and retriever models. We especially wish to thank Benedikt Schifferer, Gabriel de Souza P. Moreira, Radek Osmulski, Mengyao Xu, Ronay Ak, and Even Oldridge for providing the data from NV-Retriever (Moreira et al., 2024).
Appendix A Air Benchmark
In this first appendix section, we present AIR-Bench [^3] https://github.com/AIR-Bench/AIR-Bench (version of 24.04) that is newly released information retrieval benchmark, incorporating the diverse and comprehensive domains such as healthcare, law, news, book, arxiv, finance and synthetically generated samples using diverse LLMs. Importantly, AIR-Bench can help us to understand the generalization capability of the embedding/retrieval model, because the majority of different domain samples do not appear in MTEB benchmarks. Moreover, the AIR-Bench is designed as a closed-book benchmark whose ground truth is kept confidential. As a result, the benchmark score can be only obtained through the HuggingFace Hub platform.
In AIR-Benchmark 24.04 version, there are two tasks: QA and Long-Doc. We run evaluations on 8 English datasets in QA task and 15 English datasets on the Long-Doc task. As shown in Table 6, our NV-Embed-v2 achieves the second highest scores in QA section. As described in Table 7, our NV-Embed-v2 attained the highest scores of 74.78 on the Long-Doc section, surpassing the Bge-en-icl model that requires overheads adding in-context examples to query during training. It is important to highlight that the NV-Embed-v2 model, which achieved higher MTEB accuracy scores, also demonstrates improved accuracy on both QA and Long-Doc tasks in the AIR-Bench compared to NV-Embed-v1. Interestingly, this is not always observed in the literature, where a model performing better on MTEB does not necessarily outperform on the AIR-Bench. For example, while SFR-Embedding-2R substantially outperforms SFR-Embedding-Mistral in MTEB scores (SFR-Embedding-2R: 70.31, SFR-Embedding-Mistral: 67.56), it falls short in AIR-Bench performance both in QA (SFR-Embedding-2R: 49.47, SFR-Embedding-Mistral: 51.58) and Long-doc (SFR-Embedding-2R: 67.45, SFR-Embedding-Mistral: 69.0).
Table 6: QA (nDCG@10 scores) on AIR benchmark 24.04
Domain Wiki Web News Healthcare Law Finance Arxiv Msmarco Avg (8)
Bge-en-icl (zero-shot) 64.61 54.40 55.11 57.25 25.10 54.81 48.46 63.71 52.93
NV-Embed-v2 65.19 52.58 53.13 59.56 25.00 53.04 48.94 60.8 52.28
SFR-Embedding-Mistral 63.46 51.27 52.21 58.76 23.27 56.94 47.75 58.99 51.58
Stella-1.5B-v5 61.99 50.88 53.87 58.81 23.22 57.26 44.81 61.38 51.53
Gte-Qwen2-7B-instruct 63.46 51.20 54.07 54.20 22.31 58.20 40.27 58.39 50.26
NV-Embed-v1 62.84 50.42 51.46 58.53 20.65 49.89 46.10 60.27 50.02
Linq-Embed-Mistral 61.04 48.41 49.44 60.18 20.34 50.04 47.56 60.50 49.69
SFR-Embedding-2R 63.72 48.77 51.14 55.86 20.98 54.78 42.84 57.66 49.47
E5-mistral-7b-instruct 61.67 44.41 48.18 56.32 19.32 54.79 44.78 59.03 48.56
Table 7: Long-document (Recall@10 scores) on AIR benchmark 24.04
Domain Arxiv (4) Book (2) Healthcare (5) Law (4) Avg. (15)
NV-Embed-v2 79.27 77.46 73.01 71.18 74.78
Bge-en-icl (zero-shot) 78.30 78.21 73.65 67.09 73.75
NV-Embed-v1 77.65 75.49 72.38 69.55 73.45
Bge-multilingual-gemma2 71.77 76.46 73.96 70.86 72.88
Linq-Embed-Mistral 75.46 73.81 71.58 68.58 72.11
Stella-1.5B-v5 73.17 74.38 70.02 69.32 71.25
SFR-Embedding-Mistral 72.79 72.41 67.94 64.83 69.0
Text-embed-3-large (OpenAI) 74.53 73.16 65.83 64.47 68.77
E5-mistral-7b-instruct 72.14 72.44 68.44 62.92 68.49
SFR-Embedding-2R 70.51 70.22 67.60 62.82 67.45
Appendix B Experimental Details and Instruction Templates for Training and Evaluation
We use the Adam optimizer for each training stage. The optimizer hyperparameters are included in Table 8. We restart the optimizer with the same 50 warm-up steps and lower learning rate for the second stage.